OpenStudy (anonymous):

Solve the triangle. B = 36°, a = 41, c = 20

OpenStudy (anonymous):

b ≈ 27.5, C ≈ 115.5, A ≈ 28.5 b ≈ 45.3, C ≈ 25.5, A ≈ 118.5 b ≈ 45.3, C ≈ 29.5, A ≈ 114.5 b ≈ 27.5, C ≈ 25.5, A ≈ 118.5

terenzreignz (terenzreignz):

Law of Cosines?

geerky42 (geerky42):

Yes.

terenzreignz (terenzreignz):

You arrived at a solution, @TNNG ?

geerky42 (geerky42):

Use law of cosines to find b then use law of sines to find the rest of angles.

OpenStudy (anonymous):

ok so just use to law of cosines and i should get the answer

geerky42 (geerky42):

Well, you're not finished, since you still have to find A and C so use law of sines for this.

OpenStudy (anonymous):

i don't get it i got the D for my answer is that correct

geerky42 (geerky42):

What don't you get?

OpenStudy (anonymous):

the law of sines

geerky42 (geerky42):

The Law of Sines: \(\Large \dfrac{\sin A}{a} =\dfrac{\sin B}{b} =\dfrac{\sin C}{c}\)

geerky42 (geerky42):

Familiar with it?

OpenStudy (anonymous):

i know that but i don't get how to plug it is plus i tried it again and got Choice B

terenzreignz (terenzreignz):

Well, you could certainly use the law of Cosines to arrive at all the other missing parts of the triangle, but that's really cumbersome, compared to the simpler looking law of sines.

geerky42 (geerky42):

Yeah, it's only way to find the b. You just have to use it once, then use law of sines and you're good to go. You don't have to use entire formula, just use part of it, like \(\dfrac{\sin B}{b} = \dfrac{\sin C}{c}\) Is this clear?

geerky42 (geerky42):

The law of tangents is really cumbersome compared to the law of cosines :P

OpenStudy (anonymous):

ok so i know b=27.5

geerky42 (geerky42):

Yes.

OpenStudy (anonymous):

so figure out that a = 28.5

geerky42 (geerky42):

No, no, you already have the given value of a, so just use law of sines to find the angle.

geerky42 (geerky42):

Now you know a, b, c, and B. Now all you need to do is find A and C.

OpenStudy (anonymous):

is the angle 25.5

geerky42 (geerky42):

None, not what I get.

OpenStudy (anonymous):

115.5

pooja195 (pooja195):

b^2 = a^2 + c^2 - 2accosB b = 27.5

geerky42 (geerky42):

Wait, you're right, angle C is close to 25.5º. My bad. Now find the other angle.

geerky42 (geerky42):

180º - (A + C) = B

OpenStudy (anonymous):

118.5

OpenStudy (anonymous):

so i had it right the first time

geerky42 (geerky42):

Yes! Is this clear?

geerky42 (geerky42):

You have to do math, though. You can't just guess. Just saying.